# Eirini Schoinas1, Adyah Rastogi2, Anissa Carter3, Jacob Granley2, and Michael Beyeler2,4

### v1

@@ I Introduction
To address these nonlinearities, computational models have been developed to predict perceptual responses to electrical stimulation. These forward models use user-specific parameters to capture how stimulus properties affect the brightness and shape of phosphenes, as well as their interaction across electrodes [15, 9, 16, 17, 18]. While forward models enable a mechanistic understanding of prosthetic vision, their utility in real-world applications hinges on the ability to invert them: that is, to determine the optimal stimulus parameters required to elicit a desired percept. Deep stimulus encoders (DSEs) have been proposed for this purpose (Fig. 1A), leveraging deep learning to approximate the inverse mapping from percepts to stimuli [19, 20, 21, 17, 22]. However, DSEs require precise knowledge of user-specific parameters, which is often infeasible due to limited data availability or the inherent variability across users [23].
To optimize stimulus strategies from limited human feedback, the human-in-the-loop optimization (HILO) framework was introduced by Tristan and Fauvel [24] (Fig. 1B). Their work applied HILO to a linear phosphene model, demonstrating the feasibility of estimating parameters using pairwise comparisons of two competing simulated percepts (duels). Granley et al. [23] extended this approach by incorporating DSEs and testing it under more realistic conditions, such as model misspecifications and noisy user feedback. Their results suggested that HILO could efficiently adapt stimulus encoders to individual users, even when users selected duel outcomes randomly in two out of three duels. However, this work relied entirely on simulated users whose perceptual decisions followed predefined rules based on model predictions. Human decision making, by contrast, is inherently more variable and may not conform to model assumptions, leaving open the question of whether HILO would perform similarly when real users provide feedback.
@@
In this study, we take an intermediate step towards real-world validation by incorporating sighted participants viewing simulated prosthetic vision (SPV) stimuli on a monitor (Fig. 1C). This approach allows us to systematically test whether HILO can effectively adapt to individual perceptual variability, maintain robustness to model misspecifications, and remain practical under real-world decision-making constraints. By replacing simulated patient decisions with human feedback, this work provides critical evidence for the utility of HILO in personalizing neuroprosthetic devices.
@@ II-B Task
Participants viewed simulated prosthetic vision on a monitor and completed a series of trials aimed at optimizing the mapping between electrical stimuli and perceived visual representations. Each trial, referred to as a “duel,” presented two phosphenes generated for a target image. Participants selected the phosphene they perceived to better match the target image, described textually (e.g., “number eight”) rather than visually, to encourage independent mental representations of the target. To prevent selection bias, the positions of the two stimuli (left or right) were randomized.
In practice, electrical stimulation can evoke phosphenes that are excessively bright or even painful. To account for the limited dynamic range of computer monitors, numerical brightness values were displayed alongside each stimulus. During the tutorial phase, participants were introduced to the brightness scale: 1 represented threshold brightness (too dim), 2 was twice the threshold (ideal), and higher values indicated increasing excessive brightness (Fig. 2).
@@
  • Optimization phase: Participants completed 60 duels, during which the HILO framework iteratively refined the simulated patient parameters $\phi$ based on their preferences.
@@ II-D Human-in-the-Loop Optimization (HILO)
Following [23], a neural network DSE was trained to invert the forward model, generating stimulus parameters that elicit a phosphene most closely matching a target image. The DSE was designed as a fully connected feedforward network with multiple residual blocks, each containing batch normalization and leaky ReLU activations to stabilize training and improve generalization. The model takes as input the target image and the user-specific parameters $\phi$, which encode individual variations in phosphene perception. By training the DSE across a broad distribution of $\phi$ values [23], it learns to produce optimized stimuli tailored to diverse patient profiles.
@@
where $\phi_{1}$ was the “champion” (the current best parameter set), and $\phi_{2}$ was the “challenger” (the parameter set for which preferences were most uncertain). Here, $\mathbb{E}$ and $\mathbb{V}$ denoted the expectation and variance, respectively.
@@ II-E Experimental Conditions
  • Main experiment: This was the standard experiment adapted from [23], where the unknown parameters $\phi$ for simulated patients were iteratively optimized using participant feedback.
@@
  • Out-of-distribution (OOD): In this experiment, the ground-truth $\phi$ values for simulated patients were drawn from a distribution outside the training range of the deep stimulus encoder (DSE), evaluating the framework’s ability to generalize to unseen parameter configurations.
@@ IV-A HILO Stimuli Are Preferred Across Conditions
Across all experimental conditions, participants consistently favored HILO-optimized stimuli over the baseline encoders. Even when the assumed model parameters were misspecified, HILO maintained its advantage, demonstrating robustness to variations in user-specific parameters.
@@
IV-B Human Decision-Making Differs from Simulated Patients
Previous studies evaluating HILO [24, 23] relied entirely on simulated patients, whose decisions followed predefined loss functions. While Granley et al. [23] attempted to account for human variability by introducing artificial noise (i.e., demonstrating that HILO remained effective even when simulated patients made random choices in two out of three trials), real human decision-making may not align with these assumptions.
@@ IV-C Limitations and Future Work
While this study provides strong support for the use of HILO in personalizing neuroprosthetic devices, key challenges remain. The next step is testing HILO with blind prosthesis users, but this presents practical and safety considerations. Deep learning models can generate unpredictable outputs, and while neuroprosthetic devices include firmware to enforce safety limits, additional validation is needed to ensure that optimized stimuli do not produce unintended effects. Future work should focus on refining the optimization objective to better align with human perception, possibly incorporating alternative perceptual loss functions that capture high-level visual features rather than pixel-wise accuracy.
Beyond visual prostheses, this approach may generalize to other sensory neuroprosthetic devices, including cochlear implants and tactile feedback systems, where forward models and deep encoders are already in use [27, 28, 29, 30]. By demonstrating HILO’s effectiveness in human participants, this study marks a significant step toward fully individualized neuroprosthetic optimization.

### v2

@@ Abstract
Clinical Relevance—Validating HILO with sighted participants viewing simulated prosthetic vision is an important step toward personalized calibration of future visual prostheses.
@@ I Introduction
To address these nonlinearities, computational models have been developed to predict perceptual responses to electrical stimulation. These forward models use user-specific parameters to capture how stimulus properties affect the brightness and shape of phosphenes, as well as their interaction across electrodes [15, 9, 16, 17, 18]. While forward models enable a mechanistic understanding of prosthetic vision, their utility in real-world applications hinges on the ability to invert them: that is, to determine the optimal stimulus parameters required to elicit a desired percept. Deep stimulus encoders (DSEs) have been proposed for this purpose (Fig. 1A), leveraging deep learning to approximate the inverse mapping from percepts to stimuli [19, 20, 21, 17, 22]. However, DSEs require precise knowledge of user-specific parameters, which is often infeasible due to limited data availability or the inherent variability across users [23]. Moreover, interviews with blind prosthesis users have highlighted gaps between laboratory models and real-world experiences, underscoring the importance of designing technologies that adapt to individual needs and usability constraints [24].
To optimize stimulus strategies from limited human feedback, the human-in-the-loop optimization (HILO) framework was introduced to the field by Fauvel and Chalk [25] (Fig. 1B). Their work applied HILO to a linear phosphene model, showing that parameter estimation was feasible using pairwise comparisons of simulated percepts (duels). Granley et al. [23] extended this approach by incorporating DSEs and testing it under more realistic conditions, including model misspecifications and noisy user feedback. Their results suggested that HILO could effectively personalize stimulus encoding, even when simulated users selected randomly in two out of three trials. However, these simulations relied on predefined probabilistic rules, where choices were more likely when a percept had a lower computed error metric. Human decision making, by contrast, is inherently more variable and may not align with model assumptions, leaving open the question of whether HILO would perform similarly when real users provide feedback.
@@
In this study, we take an intermediate step toward real-world validation by incorporating sighted participants viewing simulated prosthetic vision (SPV) stimuli on a monitor (Fig. 1C). This approach enables systematic testing of whether HILO can adapt to individual perceptual variability, maintain robustness to model misspecifications, and remain practical under real-world decision-making constraints. Importantly, we evaluated HILO not only under standard conditions but also in out-of-distribution scenarios, testing its generalizability beyond the training range. We also systematically varied key phosphene properties, such as size and elongation, highlighting how perceptual distortions affect optimization performance and motivating future efforts to improve phosphene focality. By replacing simulated user decisions with human feedback, this study provides critical evidence for the utility of HILO in personalizing neuroprosthetic devices.
@@ II-B Task
Participants viewed simulated prosthetic vision on a monitor and completed a series of trials aimed at optimizing the mapping between simulated electrical stimuli and perceived visual representations. Each trial, referred to as a “duel,” presented two full percepts generated for a target image. Participants selected the percept they judgedt to better match the target, which was described textually (e.g., “number eight”) to encourage independent mental representations. To prevent selection bias, the positions of the two stimuli (left or right) were randomized across trials.
In practice, electrical stimulation can evoke phosphenes that are excessively bright or even painful [26, 27, 11, 3]. To account for the limited dynamic range of computer monitors, numerical brightness values were displayed alongside each stimulus. During the tutorial phase, participants were introduced to the brightness scale: 1 represented threshold brightness (too dim), 2 was twice the threshold (ideal), and higher values indicated increasing excessive brightness (Fig. 2).
@@
  • Optimization phase: Participants completed 60 duels, during which the HILO framework iteratively refined the simulated user parameters $\phi$ based on their preferences.
@@ II-D Human-in-the-Loop Optimization (HILO)
Following [23], a neural network DSE was trained to invert the forward model, generating stimulus parameters that elicit a phosphene most closely matching a target image. The DSE was designed as a fully connected feedforward network with multiple residual blocks, each containing batch normalization and leaky ReLU activations to stabilize training and improve generalization. The model takes as input the target image and the user-specific parameters $\phi$, which encode individual variations in phosphene perception. By training the DSE across a broad distribution of $\phi$ values [23], it learns to produce optimized stimuli tailored to diverse user profiles.
@@
where $\phi_{1}$ was the “champion” (the current best parameter set), and $\phi_{2}$ was the “challenger” (the parameter set for which preferences were most uncertain). Here, $\mathbb{E}$ and $\mathbb{V}$ denote the expectation and variance, respectively.
@@ II-E Experimental Conditions
  • Main experiment: This was the standard experiment adapted from [23], where the unknown parameters $\phi$ for simulated users were iteratively optimized using participant feedback.
@@
  • Out-of-distribution (OOD): In this experiment, the ground-truth $\phi$ values for simulated users were drawn from a distribution outside the training range of the deep stimulus encoder (DSE), evaluating the framework’s ability to generalize to unseen parameter configurations.
@@ IV-A HILO Stimuli Are Preferred Across Conditions
Across all experimental conditions, participants consistently favored HILO-optimized stimuli over baseline encoders. Even when model parameters were intentionally misspecified, HILO maintained its advantage, demonstrating robustness to variations in user-specific parameters. These results suggest that, even when the underlying model is imperfect, HILO can continue to improve performance by optimizing toward patient preferences, which remain informative despite deviations from model assumptions.
@@
IV-B Human Decision-Making Differs from Simulated Users
Previous studies evaluating HILO [25, 23] relied entirely on simulated users, whose decisions followed predefined loss functions. While Granley et al. [23] attempted to account for human variability by introducing artificial noise (i.e., demonstrating that HILO remained effective even when simulated users made random choices in two out of three trials), real human decision-making may not align with these assumptions.
@@ IV-C Limitations and Future Work
While this study provides strong support for the use of HILO in personalizing neuroprosthetic devices, several important challenges remain. The next step is testing HILO with blind prosthesis users, though this presents practical and safety considerations. Deep learning models can produce unpredictable outputs, and while most neuroprosthetic devices include firmware safeguards to enforce stimulation limits, additional validation will be needed to ensure that optimized stimuli do not generate unintended or adverse effects.
Additional limitations of the present study include the use of a static computer screen to display simulated percepts, which does not fully capture the dynamic experience of artificial vision in everyday life. Moreover, the simulated percepts shown to participants did not incorporate the temporal fading that real prosthesis users experience, potentially affecting the ecological validity of participant responses.
Future work should refine the optimization objective to better align with human perceptual judgments, moving beyond pixel-wise error metrics toward loss functions that capture higher-level visual features [21, 30]. Extending validation to blind participants and testing under more dynamic, real-world conditions will be critical for clinical translation. Recent work underscores the need to align prosthetic development with user needs [24] and the realities of assistive technology use in daily life [31], emphasizing that future optimization must prioritize not only perceptual similarity but also usability and functional task performance. Additionally, optimizing stimulation to align with latent neural representations in the visual cortex [32] offers a promising path toward deeper integration between artificial stimulation and biological processing. Together, these directions move beyond technical optimization toward meaningful, user-centered neuroprosthetic design.
Beyond visual prostheses, the HILO framework may generalize to other sensory neuroprosthetic systems, such as cochlear implants and tactile feedback devices, where forward models and deep encoders are increasingly being adopted [33, 34, 35, 36]. Recent work on brain co-processors [37] emphasizes the growing importance of closed-loop systems that jointly optimize neural decoding and stimulation in collaboration with the brain itself. Our findings support this broader vision by demonstrating that real human feedback can successfully guide the adaptation of stimulus encoding strategies, advancing toward tighter integration of artificial and biological neural systems. By establishing the viability of HILO with human participants, this study represents an important step toward fully individualized optimization of neuroprosthetic technologies.